VELOCITY MODULATED LASER ABSORPTION SPECTROSCOPY OF TiCl^{+}: ANALYSIS OF THE [17.9]^{3}Δ-(1)^{3}Δ TRANSITION

\underline{LEONID A. KALEDIN} AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The [17.9]^{3}Δ-(1)^{3}Δ (0,0) band of TiCl^{+} in the 17400-17500 cm^{−1} region has been studied using a dual-beam laser absorption technique. Spectra have been recorded using velocity modulation ^{a} at the frequency of the discharge (1f mode) and concentration modulation at the double frequency of the discharge (2f mode). TiCl^{+} ions were produced in a 1 m long, 10 mm internal diameter glass cell. The discharge was modulated at 25 kHz by a power supply which produced a bipolar sine wave. Principal molecular parameters for the (1)^{3}Δ state of the ^{48}Ti^{35}Cl^{+} isotopomer are:

T_{0}((1)^{3}Δ_{3}) = 635.191(4), B_{0} = 0.178773(8),
T_{0}((1)^{3}Δ_{2}) = 558.614(3), B_{0} = 0.179686(8),
T_{0}((1)^{3}Δ_{1}) = 431.260(2), B_{0e} = 0.180470(8), B_{0f} = 0.180453(7).

Λ-doubling was observed in the (1)^{3}Δ_{1} component of the (1)^{3}Δ state. As no (1)^{3}Π state has been observed, the parity assignment was arbitrarily chosen such that the e- parity component lies above the f- parity component.

Work supported by AFOSR and the Mission Research Corporation under grant F19628-93-C-0186.

aC. S. Gudeman, C. C. Martner, and R. J. Saykally, Phys. Rev. Lett. 50, 727 (1983).